# Bayesian Optimization with Gaussian Processes to Accelerate Stationary Point Searches

### 16 Mar 2026 v2

@@ 6 GPR-Accelerated Nudged Elastic Band
GP-NEB instantiates Algorithm 4 with NEB force relaxation as the inner optimization and explicit UCB acquisition from unevaluated images as the selection criterion. The AIE variant uses exhaustive acquisition (all $P$ images per iteration); the OIE variant selects a single image via the variance criterion (Eq. 6.1). Per-bead FPS and EMD trust adapt the shared components to the string discretization. The additional design choice specific to NEB is how many images to evaluate at each outer iteration, which determines the trade-off between surrogate accuracy and the cost per cycle [55, 53, 14, 99, 26].
@@
This active learning criterion selects the image where the surrogate is least certain, roughly halving the evaluations compared to AIE. Equation 6.1 is the pure-variance criterion ($\kappa=0$ in Eq. 4.3); the LCB-augmented form balances force magnitude against uncertainty and is the default in chemgp-core. The trust region safeguards from Section 5.2 apply to both variants. When an image drifts beyond the reliable region of the surrogate, the constraint violation triggers an evaluation at that image, which is an implicit acquisition strategy. Figure S2 (right) illustrates both variants.
@@
Figure 12: Muller-Brown potential energy surface with NEB path overlay. Filled contours show the energy landscape with three local minima (A, B, C) and two saddle points (S1, S2). Eleven NEB images (coral circles, numbered) trace the minimum energy path from A to B through S2. The climbing image (highest-energy interior image) approximates the saddle point. Energy values clipped to [200,50][-200,50] a.u.
@@ 7 GPR-Accelerated Minimization
Input: Initial configuration $\mathbf{x}$, force threshold $\epsilon_{\text{force}}$
Output: minimum at $\mathbf{x}$
1 Evaluate $V(\mathbf{x}),\mathbf{F}(\mathbf{x})$ on true PES;
2 repeat
3   Add $(\mathbf{x},V,\nabla V)$ to training set $\mathcal{D}$;
4   Re-optimize GP hyperparameters (Eq. 3.14);
5   Run L-BFGS on $V_{\text{GP}}(\mathbf{x})$ until GP forces vanish or trust radius violated;
6   Evaluate true PES at new $\mathbf{x}$;
7 if $|\mathbf{F}(\mathbf{x})|<\epsilon_{\text{force}}$ then
8    return minimum at $\mathbf{x}$;
9
10   end if
11
12until converged;
@@ 6   Evaluate true PES at new $\mathbf{x}$;
Figure 15: Convergence comparison of the GP-minimizer and classical L-BFGS on the LEPS surface. The GP surrogate reaches the convergence threshold in 10 oracle calls, compared to roughly 200 for direct gradient descent.
@@
8 The Optimal Transport Gaussian Process Framework
@@
    How do we keep the surrogate honest? The signal variance can run away and the hyperparameters can oscillate, both of which produce surrogates that are unrelated to the true PES (Section 8.2).
  2. 2.
    How far should we trust the surrogate? A fixed trust radius is either too conservative (wasting oracle calls) or too aggressive (producing unphysical steps). The threshold should reflect how much the GP has actually learned (Section 8.3).
  3. 3.
    Which training points matter? As the dataset grows, the cubic cost of hyperparameter optimization becomes the bottleneck. Selecting a geometrically diverse subset keeps the cost bounded without sacrificing surrogate quality (Section 8.1).
The generic Bayesian surrogate loop (Algorithm 4) was introduced in Section 4. The present section develops the four shared components that make that loop reliable for production use: FPS with EMD (Section 8.1), MAP regularization (Section 8.2), the adaptive trust radius (Section 8.3), and RFF scaling (Section 8.4). Table 3 summarizes how these components specialize for each method.
Table 3: Shared Bayesian optimization components across GP methods. Each column shows how the six components specialize for a given method. Acquisition modes and the oracle gate are formalized in Section 4.
Component Minimize GP-dimer OTGPD NEB AIE NEB OIE FPS subset global global HOD per-bead per-bead Trust metric Euclid./EMD EMD EMD EMD EMD RFF predict optional optional optional optional optional Acquisition implicit (Eq. 4.2) LCB + trust LCB + adaptive $T_{\text{GP}}$ exhaustive UCB (Eq. 4.3) Oracle gate $\sigma_{\perp}<\sigma_{\text{gate}}$ $\sigma_{\perp}<\sigma_{\text{gate}}$ $\sigma_{\perp}$ phase Inner optim L-BFGS rot.+trans. rot.+trans. L-BFGS L-BFGS
@@ 8.1  Farthest Point Sampling with Earth Mover’s Distance
The distance metric $D$ in Eq. 8.1 determines which configurations the algorithm considers “similar” and which it considers “different.” A bad choice here can sabotage the entire FPS selection. The 1D-max-log distance (Eq. 5.2) compares atoms by fixed index, and this breaks down whenever chemically equivalent atoms swap positions. The classic example is a methyl group rotation: three hydrogens rotate by 120 degrees, but because each hydrogen keeps its original index, the fixed-index metric registers a large geometric displacement even though the molecule has barely changed. The trust region then rejects the configuration as “too far” from the training data, or FPS treats two nearly identical structures as maximally different, wasting a slot in the subset. What we need is a distance that matches atoms of the same element before measuring displacement.
@@
The per-type linear assignment problem (Eq. 8.2) is a discrete optimal transport problem, which gives the framework its name. The “earth” being moved is a unit point mass at each atomic position, and the “work” is the total displacement needed to transform one configuration into the other [29].
@@ 9.3  PET-MAD Molecular System
Figure 23: Reaction valley projection [33] of the NEB paths on the PET-MAD surface. Reaction progress $s$ and orthogonal deviation $d$ are projected RMSD coordinates. Standard NEB (blue stars) and GP-NEB AIE (orange stars) saddle points are compared; both lie near the true saddle (black square). Dashed contours show GP variance ($\sigma^{2}$); the paths remain within the low-variance region where the surrogate is reliable. Bottom: molecular snapshots at key points along the reaction coordinate.
@@ Acknowledgements
11 Supporting Information ignoreheading:appendix
The following supplementary material provides additional details on algorithm flowcharts, marginal likelihood landscapes, hyperparameter defaults, and code implementation details.
.1  Algorithm Flowcharts
@@ 11 Supporting Information ignoreheading:appendix
Appendix A Marginal Likelihood Landscape
@@
Appendix B Hyperparameter Defaults
@@
Appendix C Connection to Code: chemgp-core
@@
C.1  Kernel Evaluation and Analytical Derivative Blocks
@@
C.2  Covariance Matrix Assembly and GP Training
@@ C.3  Code Listings
Listing S1. Derivative blocks for the inverse-distance kernel (kernel_blocks). The chain-rule structure projects the feature-space Hessian to Cartesian coordinates via the inverse-distance Jacobians.
@@
Listing S2. MAP-regularized hyperparameter training via SCG (train_model). All hyperparameters are optimized in log-space; Cholesky failure returns infinity to reject infeasible configurations.
@@
Listing S3. GP-dimer main loop (gp_dimer). The outer loop evaluates the oracle and grows the training set; the inner loop performs rotation and translation on the GP surface.
@@
Listing S4. GP-NEB OIE acquisition loop (gp_neb_oie). The select_image function supports MaxVariance, Ucb, and ExpectedImprovement strategies.
@@
Listing S5. Random Fourier feature construction (build_rff). Frequency vectors are sampled from the kernel’s spectral density (Bochner’s theorem); the constant kernel adds one extra basis function.
@@
Appendix D Mathematical Derivations
D.1  Rigid-Body Mode Basis Construction
@@
$\mathbf{t}_{x}$ $=(1,0,0,1,0,0,\ldots,1,0,0)^{T},$ (D.1)
$\mathbf{t}_{y}$ $=(0,1,0,0,1,0,\ldots,0,1,0)^{T},$ (D.2)
$\mathbf{t}_{z}$ $=(0,0,1,0,0,1,\ldots,0,0,1)^{T}.$ (D.3)
@@
$\mathbf{r}_{x}$ $=(0,z_{1}-z_{0},-(y_{1}-y_{0}),\ldots,0,z_{N}-z_{0},-(y_{N}-y_{0}))^{T},$ (D.4)
$\mathbf{r}_{y}$ $=(-(z_{1}-z_{0}),0,x_{1}-x_{0},\ldots,-(z_{N}-z_{0}),0,x_{N}-x_{0})^{T},$ (D.5)
$\mathbf{r}_{z}$ $=(y_{1}-y_{0},-(x_{1}-x_{0}),0,\ldots,y_{N}-y_{0},-(x_{N}-x_{0}),0)^{T}.$ (D.6)
@@
$\mathbf{u}_{1}$ $=\frac{\mathbf{t}_{x}}{\|\mathbf{t}_{x}\|},$ (D.7)
$\mathbf{u}_{2}$ $=\frac{\mathbf{t}_{y}-(\mathbf{t}_{y}\cdot\mathbf{u}_{1})\mathbf{u}_{1}}{\|\mathbf{t}_{y}-(\mathbf{t}_{y}\cdot\mathbf{u}_{1})\mathbf{u}_{1}\|},$ (D.8)
$\mathbf{u}_{3}$ $=\frac{\mathbf{t}_{z}-\sum_{j=1}^{2}(\mathbf{t}_{z}\cdot\mathbf{u}_{j})\mathbf{u}_{j}}{\|\mathbf{t}_{z}-\sum_{j=1}^{2}(\mathbf{t}_{z}\cdot\mathbf{u}_{j})\mathbf{u}_{j}\|},$ (D.9)
$\mathbf{u}_{k+3}$ $=\frac{\mathbf{r}_{k}-\sum_{j=1}^{k+2}(\mathbf{r}_{k}\cdot\mathbf{u}_{j})\mathbf{u}_{j}}{\|\mathbf{r}_{k}-\sum_{j=1}^{k+2}(\mathbf{r}_{k}\cdot\mathbf{u}_{j})\mathbf{u}_{j}\|},\quad k\in\{x,y,z\}.$ (D.10)
@@
D.2  Hyperparameter Oscillation Detection
@@
$O_{j}(t)=\begin{cases}1,&\text{if }(\theta_{j}(t)-\theta_{j}(t-1))(\theta_{j}(t-1)-\theta_{j}(t-2))<0,\\ 0,&\text{otherwise}.\end{cases}$ (D.11)
@@
D.3  Kernel Block Structure
@@
$k_{ee}$ $=k(\phi(\mathbf{x}_{1}),\phi(\mathbf{x}_{2}))$ (D.13)
$\ k_{ef}$ $=\frac{\partial k}{\partial\mathbf{x}_{2}}=\left(\frac{\partial k}{\partial\boldsymbol{\phi}}\right)^{T}\frac{\partial\boldsymbol{\phi}}{\partial\mathbf{x}_{2}}=\mathbf{J}_{2}^{T}\frac{\partial k}{\partial\boldsymbol{\phi}}$ (D.14)
$\ k_{fe}$ $=\frac{\partial k}{\partial\mathbf{x}_{1}}=\mathbf{J}_{1}^{T}\frac{\partial k}{\partial\boldsymbol{\phi}}$ (D.15)
$\ k_{ff}$ $=\frac{\partial^{2}k}{\partial\mathbf{x}_{1}\partial\mathbf{x}_{2}^{T}}=\mathbf{J}_{1}^{T}\left(\frac{\partial^{2}k}{\partial\boldsymbol{\phi}\partial\boldsymbol{\phi}^{T}}-2\frac{\partial k}{\partial\boldsymbol{\phi}}\otimes\frac{\partial\boldsymbol{\phi}}{\partial\mathbf{x}_{2}}\right)\mathbf{J}_{2}$ (D.16)
@@
D.4  Earth Mover’s Distance for Trust Regions
@@
$d_{\text{EMD}}^{(t)}=\min_{\pi\in\Pi_{t}}\sum_{i\in\mathcal{I}_{t}^{(1)}}\sum_{j\in\mathcal{I}_{t}^{(2)}}c_{ij}^{(t)}\pi_{ij}$ (D.17)
@@
$d_{\text{EMD}}(\mathbf{x}_{1},\mathbf{x}_{2})=\sum_{t}w_{t}\cdot d_{\text{EMD}}^{(t)}$ (D.18)
@@
D.5  OIE Acquisition Criterion Derivation
@@
$\alpha(\mathbf{R}_{i})=|\mathbf{F}_{i}^{\text{NEB}}|+\kappa\cdot\sigma_{\perp}(\mathbf{R}_{i},\boldsymbol{\tau}_{i})$ (D.19)

### 16 Apr 2026 v3

@@ 6 GPR-Accelerated Nudged Elastic Band
GP-NEB instantiates Algorithm 4 with NEB force relaxation as the inner optimization and explicit UCB acquisition from unevaluated images as the selection criterion. The AIE variant uses exhaustive acquisition (all $P$ images per iteration); the OIE variant selects a single image via the variance criterion (Eq. 6.1). Per-bead FPS and EMD trust adapt the shared components to the string discretization. The additional design choice specific to NEB is how many images to evaluate at each outer iteration, which determines the trade-off between surrogate accuracy and the cost per cycle [56, 54, 15, 102, 27, 81, 24].
@@
This active learning criterion selects the image where the surrogate has seen the least data in the kernel geometry (the largest posterior variance, which again is a sampling-density signal rather than a predicted error), roughly halving the evaluations compared to AIE. Equation 6.1 is the pure-variance criterion ($\kappa=0$ in Eq. 4.3); the LCB-augmented form balances force magnitude against uncertainty and is the default in chemgp-core. The trust region safeguards from Section 5.2 apply to both variants. When an image drifts beyond the reliable region of the surrogate, the constraint violation triggers an evaluation at that image, which is an implicit acquisition strategy. Figure S2 (right) illustrates both variants.
@@
Figure 12: Muller-Brown potential energy surface with NEB path overlay. Filled contours show the energy landscape with three local minima (A, B, C) and two saddle points (S1, S2). Eleven NEB images (coral circles, numbered) trace the minimum energy path from A to B through S2. The climbing image (highest-energy interior image) approximates the saddle point. Energy values are reported in the conventional Muller-Brown reduced units and clipped to the range [200,50][-200,50] for visualization.
@@ 7 GPR-Accelerated Minimization
1:Initial configuration $\mathbf{x}$, force threshold $\epsilon_{\text{force}}$
2:Evaluate $V(\mathbf{x}),\mathbf{F}(\mathbf{x})$ on true PES
3:repeat
4:  Add $(\mathbf{x},V,\nabla V)$ to training set $\mathcal{D}$
5:  Re-optimize GP hyperparameters (Eq. 3.15)
6:  Run L-BFGS on $V_{\text{GP}}(\mathbf{x})$ until GP forces vanish or trust radius violated
7:  Evaluate true PES at new $\mathbf{x}$
8:  if $|\mathbf{F}(\mathbf{x})|<\epsilon_{\text{force}}$ then
9:   return minimum at $\mathbf{x}$
10:  end if
11:until converged
@@
GP-minimization does still use the GP posterior variance, but at the inner-loop stopping level via the LCB convergence criterion (Eq. 4.2) rather than as an oracle gate. Three considerations argue for the geometric trust radius over a variance gate in the minimization setting, and their interplay clarifies why NEB uses both mechanisms while minimization uses only one.
First, calibration. Each search in this framework builds its GP from scratch and accumulates at most a few tens of observations by the time it converges. As noted when the predictive variance was introduced (Section~3), $\sigma^{2}(\mathbf{x}_{*})$ depends only on the kernel, the training locations, and the hyperparameters, and collapses to the noise floor at every training point whether or not the mean is accurate there. For a per-search GP the length scales have been fit from the same handful of observations, so a low variance in their vicinity only says that the surrogate is confident it can interpolate its own data; it is not a statement about the accuracy of the posterior mean against the true PES. Using this number as an oracle gate on each proposal would therefore over-trust the surrogate exactly when the dataset is sparse and the bias is largest. The LCB convergence criterion is less aggressive in this respect because it enters only after the inner optimizer has converged on the current surrogate; the oracle is then queried at the candidate regardless of the variance, and the variance merely tightens the inner stopping rule.
Second, cost. Each variance evaluation in the full augmented GP requires a triangular solve against the Cholesky factor of $\mathbf{K}$ whose cost is $\mathcal{O}(M^{2})$ per query and grows quadratically with the training set (Section 8.5). A variance gate that fires once per proposal adds one such solve to every inner-loop step, which in practice doubles or triples the GP overhead per outer iteration. For minimization the inner loop can run hundreds of surrogate steps, so the gate would be evaluated hundreds of times for a single oracle call that in the trust-region design costs nothing beyond a comparison.
Third, chemical relevance. The geometric trust radius in EMD (Section 5.2) measures per-atom displacement in Angstroms and is bounded below by a physical length scale ($\sim 0.3$ Å in chemgp-core) that reflects what a bond can reasonably do in one step regardless of what the surrogate reports. The variance gate has no such floor: as the surrogate fills in, $\sigma_{\perp}$ keeps shrinking even in regions where a single-step move would break physical bonds. Tying the step to displacement rather than to kernel-space distance therefore encodes the same prior that a chemist would apply by hand and costs a single inner product.
NEB sits in a different trade-off because the acquisition decision is which of $P$ path images to evaluate next, not whether to evaluate the proposed image at all. The per-image variance must be computed for the selection criterion anyway (Eq. 6.1), so the cost objection disappears and the calibration concern is mitigated by the fact that the NEB ensemble of images provides richer training data than a single-point optimizer accumulates. NEB therefore benefits from both the geometric trust region (per image, to keep the chain well-behaved) and the variance-based image selection (per outer iteration, for acquisition), while minimization benefits only from the former.
Figure 15: Convergence comparison of the GP-minimizer and classical L-BFGS on the LEPS surface. With a force convergence threshold of $10^{-2}$ eV/Å, the GP surrogate reaches the threshold in 9 oracle calls, compared with 57 for direct L-BFGS on the same starting configuration. Force values plotted on a logarithmic scale.
@@
8 Practical Components for the Bayesian Surrogate Loop
The generic Bayesian surrogate loop of Section 4 (Algorithm 4) needs four supporting components to be useful in production: a way to select training subsets so that hyperparameter optimization stays bounded, a way to keep MAP-NLL hyperparameter optimization from drifting into pathological regions, an adaptive trust radius that reflects how much the surrogate has actually learned, and a numerically stable solver for the kernel linear system. Three of these (training-subset selection, MAP regularization, adaptive trust radius) admit a core formulation that every method in this review uses, plus an OT-GP extension that further improves accuracy or stability for harder problems. The fourth component, the linear-algebra solver, is purely an implementation concern and lives in Section 8.5 alongside its OT-GP-flavored adaptive jitter strategy. Random Fourier features (Section 8.4) are a separate optional extension that scales prediction to large training sets.
Each of the following subsections leads with the core formulation and then flags the OT-GP refinement explicitly. The section answers three motivating questions, in order:
@@ 8 Practical Components for the Bayesian Surrogate Loop
    /How do we keep the surrogate honest?
/ The signal variance can run away and the hyperparameters can oscillate, both of which produce surrogates that are unrelated to the true PES (Section 8.2).
  1. 1.
    /How far should we trust the surrogate?
/ A fixed trust radius is either too conservative (wasting oracle calls) or too aggressive (producing unphysical steps). The threshold should reflect how much the GP has actually learned (Section 8.3).
  1. 1.
    /Which training points matter?
/ As the dataset grows, the cubic cost of hyperparameter optimization becomes the bottleneck. Selecting a geometrically diverse subset keeps the cost bounded without sacrificing surrogate quality (Section 8.1).
Table 3 summarizes how these components specialize for each method.
Table 3: Shared Bayesian optimization components across GP methods. Each column shows how the six components specialize for a given method. Acquisition modes and the oracle gate are formalized in Section 4; the LCB convergence criterion (Eq. 4.2) governs the inner-loop stopping rule and is distinct from the acquisition step listed here.
Component Minimize GP-dimer OTGPD NEB AIE NEB OIE FPS subset global global HOD per-bead per-bead Trust metric Euclid./EMD EMD EMD EMD EMD RFF predict optional optional optional optional optional Acquisition implicit (trust-clipped step) implicit (trust-clipped step) implicit (trust-clipped step, adaptive $T_{\text{GP}}$) exhaustive UCB (Eq. 4.3) Oracle gate $\sigma_{\perp}<\sigma_{\text{gate}}$ $\sigma_{\perp}<\sigma_{\text{gate}}$ $\sigma_{\perp}$ phase Inner optim L-BFGS rot.+trans. rot.+trans. L-BFGS L-BFGS
@@ 8.1  Farthest Point Sampling with Earth Mover’s Distance
The distance metric $D$ in Eq. 8.1 determines which configurations the algorithm considers "similar" and which it considers "different." A bad choice here can sabotage the entire FPS selection. The 1D-max-log distance (Eq. 5.2) compares atoms by fixed index, and this breaks down whenever chemically equivalent atoms swap positions. The classic example is a methyl group rotation: three hydrogens rotate by 120 degrees, but because each hydrogen keeps its original index, the fixed-index metric registers a large geometric displacement even though the molecule has barely changed. The trust region then rejects the configuration as "too far" from the training data, or FPS treats two nearly identical structures as maximally different, wasting a slot in the subset. What we need is a distance that matches atoms of the same element before measuring displacement.
@@
The per-type linear assignment problem (Eq. 8.2) is a discrete optimal transport problem, which gives the framework its name. The "earth" being moved is a unit point mass at each atomic position, and the "work" is the total displacement needed to transform one configuration into the other [30].
@@ 8.4  Scaling via Random Fourier Features
RFF fitting and prediction operate in the inverse-distance feature space $\boldsymbol{\phi}(\mathbf{x})$, not in Cartesian coordinates. The random frequencies $\boldsymbol{\omega}_{m}$ come from the spectral density of the SE kernel in feature space (Eq. 8.8), and the design matrix $\mathbf{Z}$ evaluates the cosine features at $\boldsymbol{\phi}(\mathbf{x})$ rather than at $\mathbf{x}$. The SE kernel remains stationary in $\boldsymbol{\phi}$, but the composite kernel $k(\boldsymbol{\phi}(\mathbf{x}),\boldsymbol{\phi}(\mathbf{x}^{\prime}))$ loses Cartesian stationarity because $\boldsymbol{\phi}$ depends nonlinearly on $\mathbf{x}$. Bochner’s theorem still applies in the feature space where stationarity holds, and Cartesian gradients follow by composing the RFF gradient in feature space with the inverse-distance Jacobian (Eq. 3.14) through the chain rule, exactly as in Eq. 8.11.
@@
8.5  Implementation: Cholesky Solver and Adaptive Jitter
Every prediction and every step of MAP-NLL hyperparameter optimization requires solving the linear system
$\mathbf{K}\,\boldsymbol{\alpha}=\mathbf{y},$ (8.12)
where $\mathbf{K}$ is the augmented covariance matrix of dimension $M(1+3N)$ (energies and forces stacked as in Eq.~3.8) and $\mathbf{y}$ holds the corresponding observations. A naive matrix inverse $\boldsymbol{\alpha}=\mathbf{K}^{-1}\mathbf{y}$ is both numerically unstable and unnecessarily expensive [83]. For a symmetric positive-definite $\mathbf{K}$ the standard recipe is the Cholesky factorization
$\mathbf{K}=\mathbf{L}\mathbf{L}^{\top},\quad\mathbf{L}\text{ lower-triangular},$ (8.13)
followed by two triangular solves: forward substitution $\mathbf{L}\,\mathbf{z}=\mathbf{y}$ followed by back substitution $\mathbf{L}^{\top}\boldsymbol{\alpha}=\mathbf{z}$. The factorization itself costs $\mathcal{O}(M^{3})$ flops with a small constant; the two triangular solves are $\mathcal{O}(M^{2})$ each. The Cholesky factor $\mathbf{L}$ also gives the log determinant for free, $\log\det\mathbf{K}=2\sum_{i}\log L_{ii}$, which is the dominant term in the negative log marginal likelihood that drives hyperparameter optimization. The implementation in chemgp-core calls the faer::llt routine in the Rust covariance.rs module and reuses $\mathbf{L}$ for both the linear solve and the log-determinant computation.
Two failure modes break this recipe in practice. First, as the dataset fills in, nearby configurations contribute nearly collinear rows to $\mathbf{K}$ and the smallest eigenvalue can fall below the working precision, so that $\mathbf{K}$ is positive definite analytically but indefinite in finite arithmetic. Second, during MAP-NLL optimization the proposed hyperparameters themselves can drive $\mathbf{K}$ close to singular long before the data do, especially when $\sigma_{f}^{2}$ becomes large or a length scale collapses. The textbook remedy is to add a small multiple of the identity to the diagonal of $\mathbf{K}$: write
$\widetilde{\mathbf{K}}(\eta)=\mathbf{K}+\eta\,\mathbf{I},$ (8.14)
and choose $\eta\geq 0$ just large enough that the Cholesky factorization of $\widetilde{\mathbf{K}}(\eta)$ succeeds. The added $\eta$ acts as Tikhonov regularization on Eq.~8.12 and shifts the spectrum of $\mathbf{K}$ by $\eta$ without changing its eigenvectors.
The OT-GP refinement here is to set $\eta$ adaptively rather than to a fixed constant. The implementation starts with $\eta_{0}=10^{-8}\,\max(\mathrm{diag}(\mathbf{K}))$, scaled to the matrix so that the same code is meaningful across the eV (molecular) and reduced-unit (model surface) regimes. If the Cholesky of $\widetilde{\mathbf{K}}(\eta_{0})$ fails, the jitter is increased geometrically, $\eta_{k+1}=10\,\eta_{k}$, and the factorization is retried, up to a configurable retry limit. The first $\eta_{k}$ that succeeds is the regularizer used for that step. This gives a one-line answer to the reviewer-2 question of a lower bound on the regularizer: the bound is the smallest $\eta_{k}$ in this geometric ladder that keeps $\widetilde{\mathbf{K}}(\eta_{k})$ positive definite, which is set by the data and the current hyperparameters rather than fixed in advance. We tried fixed lower bounds and found the adaptive ladder more robust across the regimes the tutorial covers.
A complementary safeguard works at the hyperparameter optimization level rather than on $\mathbf{K}$ directly. The MAP-NLL surface develops a runaway-variance basin in which the optimizer pushes $\sigma_{f}^{2}\to\infty$ to absorb residual error; without a check, the SCG optimizer will follow it. The barrier of Eq.~8.4 caps $\log\sigma_{f}^{2}$ at $\lambda_{\max}=\ln(2)$, and the negative log-likelihood evaluation returns $\infty$ whenever this upper bound is violated. Combined with the adaptive jitter on the linear-algebra side, the GP training loop never aborts on a numerically singular $\mathbf{K}$: the solver finds a viable $\eta_{k}$, the NLL evaluation either returns a finite value or a sentinel $\infty$, and the SCG optimizer is steered back into the feasible region.
@@ 9.3  PET-MAD Molecular System
Figure 23: Reaction valley projection [34] of the NEB paths on the PET-MAD surface. Reaction progress $s$ and orthogonal deviation $d$ are projected RMSD coordinates. Standard NEB (blue stars) and GP-NEB AIE (orange stars) saddle points are compared; both lie near the true saddle (black square). Dashed contours show GP variance ($\sigma^{2}$); the paths remain within the well-sampled region where the surrogate has accumulated data, which in combination with the per-step trust region is what keeps them close to the true saddle. Bottom: molecular snapshots at key points along the reaction coordinate.
@@ 11 For Table of Contents Only
Appendix A Algorithm Flowcharts
@@
Appendix B Marginal Likelihood Landscape
@@
Appendix C Hyperparameter Defaults
@@
Appendix D Connection to Code: chemgp-core
@@
D.1  Kernel Evaluation and Analytical Derivative Blocks
@@
D.2  Covariance Matrix Assembly and GP Training
@@ D.3  Code Listings
*Listing S1.
Appendix E Derivative blocks for the inverse-distance kernel (kernel_blocks).
The chain-rule structure projects the feature-space Hessian to Cartesian coordinates via the inverse-distance Jacobians.
@@
*Listing S2.
Appendix F MAP-regularized hyperparameter training via SCG (train_model).
All hyperparameters are optimized in log-space; Cholesky failure returns infinity to reject infeasible configurations.
@@
*Listing S3.
Appendix G GP-dimer main loop (gp_dimer).
The outer loop evaluates the oracle and grows the training set; the inner loop performs rotation and translation on the GP surface.
@@
*Listing S4.
Appendix H GP-NEB OIE acquisition loop (gp_neb_oie).
The select_image function supports MaxVariance, Ucb, and ExpectedImprovement strategies.
@@
*Listing S5.
Appendix I Random Fourier feature construction (build_rff).
Frequency vectors are sampled from the kernel’s spectral density (Bochner’s theorem); the constant kernel adds one extra basis function.
@@
Appendix J Mathematical Derivations
J.1  Rigid-Body Mode Basis Construction
@@
$\mathbf{t}_{x}$ $=(1,0,0,1,0,0,\ldots,1,0,0)^{T},$ (J.1)
$\mathbf{t}_{y}$ $=(0,1,0,0,1,0,\ldots,0,1,0)^{T},$ (J.2)
$\mathbf{t}_{z}$ $=(0,0,1,0,0,1,\ldots,0,0,1)^{T}.$ (J.3)
@@
$\mathbf{r}_{x}$ $=(0,z_{1}-z_{0},-(y_{1}-y_{0}),\ldots,0,z_{N}-z_{0},-(y_{N}-y_{0}))^{T},$ (J.4)
$\mathbf{r}_{y}$ $=(-(z_{1}-z_{0}),0,x_{1}-x_{0},\ldots,-(z_{N}-z_{0}),0,x_{N}-x_{0})^{T},$ (J.5)
$\mathbf{r}_{z}$ $=(y_{1}-y_{0},-(x_{1}-x_{0}),0,\ldots,y_{N}-y_{0},-(x_{N}-x_{0}),0)^{T}.$ (J.6)
@@
$\mathbf{u}_{1}$ $=\frac{\mathbf{t}_{x}}{\|\mathbf{t}_{x}\|},$ (J.7)
$\mathbf{u}_{2}$ $=\frac{\mathbf{t}_{y}-(\mathbf{t}_{y}\cdot\mathbf{u}_{1})\mathbf{u}_{1}}{\|\mathbf{t}_{y}-(\mathbf{t}_{y}\cdot\mathbf{u}_{1})\mathbf{u}_{1}\|},$ (J.8)
$\mathbf{u}_{3}$ $=\frac{\mathbf{t}_{z}-\sum_{j=1}^{2}(\mathbf{t}_{z}\cdot\mathbf{u}_{j})\mathbf{u}_{j}}{\|\mathbf{t}_{z}-\sum_{j=1}^{2}(\mathbf{t}_{z}\cdot\mathbf{u}_{j})\mathbf{u}_{j}\|},$ (J.9)
$\mathbf{u}_{k+3}$ $=\frac{\mathbf{r}_{k}-\sum_{j=1}^{k+2}(\mathbf{r}_{k}\cdot\mathbf{u}_{j})\mathbf{u}_{j}}{\|\mathbf{r}_{k}-\sum_{j=1}^{k+2}(\mathbf{r}_{k}\cdot\mathbf{u}_{j})\mathbf{u}_{j}\|},\quad k\in\{x,y,z\}.$ (J.10)
@@
J.2  Hyperparameter Oscillation Detection
@@
$O_{j}(t)=\begin{cases}1,&\text{if }(\theta_{j}(t)-\theta_{j}(t-1))(\theta_{j}(t-1)-\theta_{j}(t-2))<0,\\ 0,&\text{otherwise}.\end{cases}$ (J.11)
@@
J.3  Kernel Block Structure
@@
$k_{ee}$ $=k(\phi(\mathbf{x}_{1}),\phi(\mathbf{x}_{2}))$ (J.13)
$\ k_{ef}$ $=\frac{\partial k}{\partial\mathbf{x}_{2}}=\left(\frac{\partial k}{\partial\boldsymbol{\phi}}\right)^{T}\frac{\partial\boldsymbol{\phi}}{\partial\mathbf{x}_{2}}=\mathbf{J}_{2}^{T}\frac{\partial k}{\partial\boldsymbol{\phi}}$ (J.14)
$\ k_{fe}$ $=\frac{\partial k}{\partial\mathbf{x}_{1}}=\mathbf{J}_{1}^{T}\frac{\partial k}{\partial\boldsymbol{\phi}}$ (J.15)
$\ k_{ff}$ $=\frac{\partial^{2}k}{\partial\mathbf{x}_{1}\partial\mathbf{x}_{2}^{T}}=\mathbf{J}_{1}^{T}\left(\frac{\partial^{2}k}{\partial\boldsymbol{\phi}\partial\boldsymbol{\phi}^{T}}-2\frac{\partial k}{\partial\boldsymbol{\phi}}\otimes\frac{\partial\boldsymbol{\phi}}{\partial\mathbf{x}_{2}}\right)\mathbf{J}_{2}$ (J.16)
@@
J.4  Earth Mover’s Distance for Trust Regions
@@
$d_{\text{EMD}}^{(t)}=\min_{\pi\in\Pi_{t}}\sum_{i\in\mathcal{I}_{t}^{(1)}}\sum_{j\in\mathcal{I}_{t}^{(2)}}c_{ij}^{(t)}\pi_{ij}$ (J.17)
@@
$d_{\text{EMD}}(\mathbf{x}_{1},\mathbf{x}_{2})=\sum_{t}w_{t}\cdot d_{\text{EMD}}^{(t)}$ (J.18)
@@
J.5  OIE Acquisition Criterion Derivation
@@
$\alpha(\mathbf{R}_{i})=|\mathbf{F}_{i}^{\text{NEB}}|+\kappa\cdot\sigma_{\perp}(\mathbf{R}_{i},\boldsymbol{\tau}_{i})$ (J.19)